OpenStudy (anonymous):

solve log(base 3)3 + log(base3)x = 5

OpenStudy (anonymous):

x = e^5 /3

OpenStudy (anonymous):

log(base 3) 3 = 1...so log(base 3)x=5-1=4 so x=3^4=81